#1 Medical X-Rays are produced by accelerating electrons through a potential difference of 25000 V toward...

#1

Medical X-Rays are produced by accelerating electrons through a potential difference of 25000 V toward a metallic target. Calculate the relativistic velocity of such X-ray producing electrons.
Answer #1

According to the special relativity,relativistic kinetic energy of an electron accelerated by V volts is given by

( - 1 ) mc2 = eV

where m - mass of the electron= 9.1 x 10-31kg

c - velocity of light = 3 x 108m/s

e - charge of the electron = 1.6 x 10-19 C

V - potential difference = 25000 V

= (1- v2/c2)-1/2; v - velocity of the electron

substituting we get

( - 1) x9.1 x 10-31x( 3x108)2 = 1.6 x 10-19 x25000

( - 1) x 8.19 x 10-14 = 4 x 10-15

( - 1) = 4 x 10-15 / 8.19 x 10-14 = 0.049

= 1 + 0.049 = 1.049

(1- v2/c2)-1/2 = 1.049

(1- v2/c2)-1= (1.049)2 = 1.1

(1- v2/c2) = 1/1.1 = 0.91

v2/c2 = 1 - 0.91 = 0.09

v2= c2 x 0.09

v = 0.3 c = 0.3 x 3 x 108 = 9 x 107 m/s